# DEER: A Benchmark for Evaluating Deep Research Agents on Expert Report Generation

### v2

@@ 5.1 Implementation Details
We evaluate report quality using an LLM-as-a-judge approach across five dimensions—Request Fulfillment, Analytical Soundness, Structural Coherence, Format & Style, and Ethics Compliance. The LLM judge receives the task query, the report being evaluated, task-specific Expert Evaluation Guidance, and a fixed set of fine-grained rubric items. It assigns a 1–10 score and a brief rationale to each item and returns the results as a JSON object mapping each item to its score and rationale. In parallel, Information Integrity and Information Sufficiency are assessed by the Information Verification Module, which extracts claims from the report and verifies them against external evidence retrieved via web search, producing metrics such as Claim Factuality and Citation Support. Details are provided in Appendix G.4. Finally, we compute the overall quality score hierarchically by aggregating scores bottom-up from rubric items to higher-level categories. Details of the scoring procedure are provided in Appendix E.3.
Figure 3: Heatmap visualizations of expert report evaluation results. (a) Criteria-wise scores across detailed evaluation categories. (b) Domain-wise scores averaged over two representative tasks randomly sampled from each domain.
@@ 5.2 Baseline Models
To compare expert report generation performance, we consider four baseline families: General LLMs (fast), LLMs+Reasoning (think), LLMs+Reasoning+WebSearch (think+search), and Deep Research (deep). Each family is instantiated using multiple model families/backbones (Qwen, Gemini, Claude, and GPT), and they differ in whether they (i) use reasoning, (ii) use web search, and (iii) employ research-system orchestration. Detailed model configurations are provided in Appendix H.
@@ 6.1 Main Results
Summarizing Table 1, most models score highly on structure, style, and ethics (Structural Coherence, Format & Style, Ethics), but lag on requirement fulfillment and analytical soundness (Req. Fulfillment, Analytical Soundness). This suggests that report-writing quality is largely mature, whereas expert-level intent alignment and evidence-based justification remain limited. Across baseline families, adding reasoning in think improves overall performance over fast. Moreover, external-information metrics—Information Integrity and Information Sufficiency—see further gains with think+search and deep. An interesting finding is that reasoning models without web search (think) outperform think+search and deep on report-quality metrics excluding information-related scores. This suggests that integrating diverse external information can blur the problem definition and argument structure. Accordingly, simply adding search and external sources may not guarantee improved report-writing quality.
@@ 6.2 Fine-grained Analysis
Figure 3(a) breaks down the performance patterns of Deep Research systems on expert report generation into fine-grained dimensions. The sub-dimension scores for Request Fulfillment, Analytical Soundness, and Structural Coherence are generally low, with scope particularly low within Request Fulfillment. This indicates that Deep Research systems often fail to clearly specify the report scope (what is covered vs. excluded) and the assumptions/constraints used to develop the report. In addition, ref_amount is low in Information Sufficiency, suggesting that systems tend to rely on a small number of references rather than leveraging a broad set of sources. Figure 3(b) shows domain-level performance of Deep Research agents. For each domain, we report results on two tasks randomly selected from that domain. Performance tends to drop uniformly across agents on relatively difficult tasks, and Physics consistently records the lowest scores compared to other domains.
@@ 6.3 Correlation with Human Evaluation
To validate the proposed evaluation method and assess the contribution of each component, we compared LLM-based evaluations against human expert judgments. For each of the 45 reports, we collected two independent ratings from domain experts whose expertise aligned with the report’s topic (90 ratings total). We computed Pearson’s r, Spearman’s ρ, and LLM–human pairwise agreement for each of the five LLM-based evaluator models, and report the average of each metric across models. Additional details on the experimental setup and human evaluation protocol are provided in Appendix I.
@@ 6.4 Inter-evaluator Reliability
To verify that the proposed evaluation method yields consistent results across different LLM evaluators, we measure inter-evaluator reliability Artstein (2017); Lee et al. (2025). Specifically, we compute Krippendorff’s $\alpha$ and the intraclass correlation coefficient (ICC) using the scores from the five LLM judges described in Section 6.3.As shown in Table 3, Vanilla achieves moderate inter-evaluator reliability, but reliability drops substantially under +Dimensions. +Granular Rubrics yields a slight recovery, though still below Vanilla. +Expert Guidance attains the highest reliability across metrics. This suggests that expert guidance clarifies what to look for under each criterion, enabling more consistent judgments across different LLM judges.
@@ 6.5 Information Verification Module Evaluation
In this section, we evaluate the proposed Information Verification module by dividing it into two stages: claim extraction and claim verification.  [^2] Detailed content is covered in Appendix F.
@@ Appendix A Comparison with Deep Research Benchmarks
In this section, we compare existing deep research benchmarks with our methodology. Table 6 extends and modifies the comparison table proposed in ResearchRubrics Sharma et al. (2025), and adds new evaluation axes to summarize how DEER differs from prior work. The ResearchRubrics table uses five axes (whether rubrics are human-written, whether experts curate tasks, whether tasks are open-ended, whether non-technical domains are included, and whether an LLM-as-judge is used). On top of this, we add three axes that we consider essential from the perspective of deep research report evaluation: (1) Claim fact-check, (2) Interpretability, and (3) Systematic criteria. Systematic criteria indicate whether the evaluation criteria form a shared hierarchical ontology across tasks, grounded in external expert norms and guidelines.
@@
The first block of the table consists of benchmarks for evaluating search and browsing agents. AcademicBrowse Zhou et al. (2025), BrowseComp Wei et al. (2025a), and Mind2Web2 Gou et al. (2025) evaluate, respectively, the ability to browse an academic corpus or the open web to produce short answers to complex queries, the ability to perform agentic search across diverse websites, and the consistency between generated answers and cited sources. These benchmarks focus on “how well the system finds information,” i.e., search and browsing strategies, and thus are one step removed from the deep research report quality evaluation that we study.
@@
The second block covers benchmarks that focus on specific component abilities or conceptualizations of deep research. ResearchBench Liu et al. (2025) and ResearcherBench Xu et al. (2025) evaluate, respectively, the ability to generate inspiration and hypotheses from papers, and to provide long-form advice on frontier AI research questions based on expert rubrics. Characterizing Deep Research Java et al. (2025) and LiveDRBench evaluate the recovery of correct claims in search tasks that require many information units and non-trivial reasoning. SPOT Son et al. (2025) measures how well a system can detect critical errors in published papers, and ExpertLongBench Ruan et al. (2025) evaluates the ability to generate long-form expert text without external search. These benchmarks finely probe partial abilities such as hypothesis generation, claim discovery, error detection, and expert writing, but do not directly evaluate the quality of full reports produced by web- or literature-based deep research agents.
The third block targets benchmarks that evaluate the quality of deep research reports themselves. ReportBench Li et al. (2025a) and DeepScholar-Bench Patel et al. (2025) measure report quality in academic survey and related-work settings using citation coverage, overlap with gold surveys, and verifiability of citations. DeepResearchBench Du et al. (2025) and LiveResearchBench Wang et al. (2025) evaluate reports for multi-domain web-based deep research tasks using LLM-generated evaluation criteria (RACE), citation-based fact-checking (FACT), and dimensions such as coverage, depth, presentation, and citation accuracy. DeepResearchArena Wan et al. (2025) automatically constructs large numbers of research tasks from seminar transcripts, and measures factual alignment and overall quality via coverage/conflict/omission (KSR/KCR/KOR) against evidence keypoints extracted from cited URLs and LLM-generated task-specific rubrics (ACE). DeepResearchGym Coelho et al. (2025) provides an offline search environment over a public web corpus and an LLM-as-judge protocol, offering a sandbox for evaluating deep research agents’ reports with respect to alignment with information needs, search faithfulness, and overall quality. ResearchRubrics Sharma et al. (2025) provides 101 deep research prompts and over 2,500 human-written rubrics, and is designed so that an LLM-as-judge assigns per-item scores by following these detailed human criteria.
Among these, we analyze DeepResearchBench, LiveResearchBench, and ResearchRubrics in more detail, as they are closest to DEER in terms of evaluation setting. DeepResearchBench and LiveResearchBench are most similar to our web-based deep research report evaluation setting, but both rely heavily on LLMs to define evaluation criteria and assign scores, and their fact-checking is applied only to a limited subset of claims with explicit citations. ResearchRubrics has the advantage that domain experts write 20–40 rubric items for each deep research prompt, enabling fine-grained capture of what constitutes a “good” answer, but it does not perform source-level claim fact-checking, and the rubric structure differs across prompts, making it challenging to interpret system behavior consistently along shared dimensions and rubrics. In contrast, DEER extracts, as exhaustively as possible, all explicit and implicit claims in a report that can be grounded in evidence documents and performs claim-level fact-checking over this set, and evaluates reports using a shared ontology of 7 high-level areas, 25 criteria, and 101 rubric items derived from expert report evaluation norms and guidelines, augmented with task-specific expert guidance. This allows DEER to avoid designing a new rubric from scratch for each task while still applying a dense set of rubric items to each report, and to interpret system performance along standard dimensions and rubric items across tasks.
@@ B.1 Topic Domain Analysis
To construct deep research tasks, we analyzed 5,842 in-house user queries collected from our deep research system to estimate real-world domain demand and to guide the benchmark’s target domain distribution. Figure 4 shows the resulting distribution. Based on this analysis, we finalized 11 topic domains: the 10 most frequent domains and an Others category that aggregates all remaining domains.
@@ B.2 HLE Subject Mapping
Accordingly, we used Humanity’s Last Exam (HLE) Phan et al. (2025) as a source of expert-written, high-difficulty seed items that align with expert-level report generation. To preserve the 11-topic domain composition derived from actual deep research logs, we mapped each domain to one of HLE’s 13 subject domains and sampled HLE items from the corresponding subjects as domain-specific seeds. At this time, we filtered candidates via a preliminary performance evaluation, excluding items that were excessively difficult for LLM-based evaluation and retaining only those within an appropriate difficulty range. As a result, we selected a total of 50 seed items across HLE’s 13 subject domains. The correspondence between deep research topic domains and HLE subject domains is summarized in Table 8.
@@ B.3 Conversion from HLE QA to Deep Research Reports
Each task query was drafted by one domain expert and cross-reviewed by another expert from the same field. Cross-review was repeated in multiple rounds as needed to check whether the reformulated task (i) is not overly narrow, (ii) requires expert-level domain expertise, (iii) specifies a sufficiently concrete research scope and direction, and (iv) does not contain excessive hints that could steer the model toward the answer or conclusion; the task was revised accordingly. Experts were individuals with a master’s degree in the relevant field or equivalent domain expertise. An example research task is provided in Box 5.
@@ B.4 Construction of Expert Evaluation Guidance
For all 50 tasks, Expert Evaluation Guidance was written in the same expert workflow simultaneously with query reformulation, so that each task’s requirements and evaluation criteria are mutually aligned. As in query reformulation, each Guidance was drafted by one domain expert and cross-reviewed by another expert from the same field. During cross-review, we checked and revised whether (i) the topic’s mandatory content elements were included at a sufficiently detailed level and written in an evaluable way, (ii) optional content or stylistic preferences were not included as mandatory evaluation elements, and (iii) requirements specified in the task prompt’s writing requirements (writing-direction instructions/additional requested constraints) were reflected without omission in the Guidance. An example Expert Evaluation Guidance is provided in Box 6.
@@ C.2 Construction Procedure and Evidence Mapping
Appendix D Foundations of Deep Research Report Evaluation Taxonomy
@@
Completeness aligns with guideline-level completeness checks that evaluate whether all required components of a report are present according to prescribed inclusion mandates across authoritative standards Bastian and Moher (2021); Boutron et al. (2010); International Organization for Standardization (2011); What Works Clearinghouse (2022); Percie du Sert et al. (2020); Global Reporting Initiative (2023); von Elm and others (2007). Scope reflects boundary-setting requirements that assess whether a report explicitly specifies its operative limits—such as eligibility conditions, assumptions, and constraints—within the structural fields defined in major evaluative frameworks Bastian and Moher (2021); National Institute of Standards and Technology (2023); American Historical Association (2024); American Economic Association (2024); Center for Open Science (2024); NASA (2016); U.S. Securities and Exchange Commission (2024); Chan and others (2013). Helpfulness corresponds to coherence and utility standards that evaluate whether the report’s core conclusions not only align logically with the evidence base but also deliver sufficient specificity and practical feasibility to comprehensively address the user’s inquiry Guyatt et al. (2013); American Psychological Association (2025); Chang et al. (2024); Institute of Education Sciences (2022); OECD (2020).
@@
Quantification reflects evaluative requirements that check whether quantitative statements in a report correspond to verifiable computations, prespecified analytic procedures, and reproducible numerical evidence as established in major methodological frameworks Guyatt et al. (2013); Joint Committee for Guides in Metrology (2008); Association for Computing Machinery (2025); International Union of Pure and Applied Chemistry (2007); International Union of Pure and Applied Physics (2010); The Econometric Society (2024); American Educational Research Association et al. (2014); Mohr and others (2024); OECD (2011); International Accounting Standards Board (2024); International Organization for Standardization (2018); Intergovernmental Panel on Climate Change (2019). Reasoning aligns with reasoning-assessment criteria that evaluate whether inferential steps are explicitly grounded in evidence, free of unsupported leaps, and consistent with theoretical or mathematical rigor European Mathematical Society (2025); American Philosophical Association (2024); American Educational Research Association (2006); American Mathematical Society (2022); U.S. Congress (2024); Tong et al. (2007).
@@
Introduction corresponds to structural-orientation requirements that assess whether a report’s opening section presents its scope and analytic pathway in accordance with prescribed organizational fields in established reporting and specification frameworks Bastian and Moher (2021); Boutron et al. (2010); IEEE Computer Society (2018); IFRS Foundation (2021); Gagnier and others (2013). Body aligns with structural-progression criteria that evaluate whether the main sections develop the promised analytical sequence without omission or drift relative to the ordered components defined in recognized guideline structures Bastian and Moher (2021); Boutron et al. (2010); IEEE Computer Society (2018); Eaton and others (2024); Gagnier and others (2013). Conclusion reflects closure-consistency requirements that examine whether final statements integrate preceding evidence without introducing unsupported expansions, consistent with the conclusion-governance rules embedded in major evaluative standards Bastian and Moher (2021); Boutron et al. (2010); Gagnier and others (2013). Section-Level corresponds to intra- and inter-section coherence checks that assess alignment, ordering, and non-duplication of content in accordance with structured specification and reporting templates in authoritative frameworks Bastian and Moher (2021); IEEE Computer Society (2018); IFRS Foundation (2021).
@@
Report Format corresponds to format-governance requirements that evaluate whether a report adheres to prescribed structural conventions for scientific communication as codified in established editorial and reporting standards Bastian and Moher (2021); Boutron et al. (2010); International Committee of Medical Journal Editors (2025); Comrie, B. and Haspelmath, M. and Bickel, B. (2015); Society for Industrial and Applied Mathematics (2024); International Phonetic Association (1999); American Political Science Association (2018); American Sociological Association (2022). Writing Quality aligns with language-precision criteria that assess clarity, specificity, and terminological consistency according to recognized guidelines for accurate and unbiased scholarly expression Boutron et al. (2010); International Committee of Medical Journal Editors (2025); University of Chicago Press (2024); Garner (2019, 2013). Paragraph Quality reflects cohesion-assessment rules that examine whether individual paragraphs follow a coherent internal logic—anchoring topic statements, supporting evidence, and transitional structure—in line with authoritative reporting and specification frameworks Bastian and Moher (2021); IEEE Computer Society (2018); Harvard Law Review Association (2020); American Sociological Association (2022). Readability corresponds to accessibility-oriented requirements that evaluate whether the narrative facilitates comprehension through appropriate signaling, explanatory devices, and presentation practices as outlined in major editorial and reporting guidelines Bastian and Moher (2021); International Committee of Medical Journal Editors (2025); Garner (2013); Global Reporting Initiative (2023).

### v3

@@ 5.1 Implementation Details
We evaluate report quality using an LLM-as-a-judge approach across five dimensions—Request Fulfillment, Analytical Soundness, Structural Coherence, Format & Style, and Ethics Compliance. The LLM judge receives the task query, the report being evaluated, task-specific Expert Evaluation Guidance, and a fixed set of fine-grained rubric items. It assigns a 1–10 score and a brief rationale to each item and returns the results as a JSON object mapping each item to its score and rationale. In parallel, Information Integrity and Information Sufficiency are assessed by the Information Verification Module, which extracts and type-classifies claims (Types A–F) and verifies the verifiable subset (Types A–C) against evidence documents retrieved from the report‘s cited sources (URLs), with semantic back-tracking to recover omitted citations for implicit claims. The module outputs quantitative metrics aligned with the Integrity/Sufficiency subdimensions (e.g., Claim Factuality, Citation Support, Evidence Coverage, Reference Reliability/Diversity), and hierarchically aggregates them into the two dimension-level scores (Appendix D.4). We report results using GPT-5.2 for Report Quality Assessment and GPT-5-mini for the Information Verification module, with an average evaluation cost of approximately \$0.5–\$1.0 per report.
Figure 3: Heatmap visualizations of expert report evaluation results. (a) Criteria-wise scores across detailed evaluation categories. (b) Domain-wise scores averaged from each domain.
@@ 5.2 Baseline Models
To compare expert report generation performance, we consider four baseline families: General LLMs (fast), LLMs+Reasoning (think), LLMs+Reasoning+WebSearch (think+search), and Deep Research (deep). Each family is instantiated using multiple model families/backbones—Qwen, Gemini, Claude, and GPT (Yang et al., 2025; Google, 2025; Anthropic, 2025; OpenAI, 2025a), and they differ in whether they (i) use reasoning, (ii) use web search, and (iii) employ research-system orchestration. Detailed model configurations are provided in Appendix G.
@@ 6.1 Main Results
Summarizing Table 1, most models score highly on structure, style, and ethics (Structural Coherence, Format & Style, Ethics), but lag on requirement fulfillment and analytical soundness (Req. Fulfillment, Analytical Soundness). This suggests that report-writing quality is largely mature, whereas expert-level intent alignment and reasoning completeness remain limited. Across baseline families, adding reasoning in think improves overall performance over fast. Moreover, external-information metrics—Information Integrity and Information Sufficiency—see further gains with think+search and deep. An interesting finding is that reasoning models without web search (think) outperform think+search and deep on report-quality metrics excluding information-related scores. This suggests that integrating diverse external information can blur the problem definition and argument structure. Accordingly, simply adding search and external sources may not guarantee improved report-writing quality.
@@ 6.2 Fine-grained Analysis
Figure 3(a) breaks down the performance patterns of Deep Research systems on expert report generation into fine-grained dimensions. The sub-dimension scores for Request Fulfillment, Analytical Soundness, and Structural Coherence are generally low, with scope particularly low within Request Fulfillment. This indicates that Deep Research systems often fail to clearly specify the report scope (what is covered vs. excluded) and the assumptions/constraints used to develop the report. In addition, ref_amount is low in Information Sufficiency, suggesting that systems tend to rely on a small number of references rather than leveraging a broad set of sources. Figure 3(b) shows domain-level performance of Deep Research agents. We observe that performance varies substantially across disciplines: agents generally achieve higher success in Philosophy, Psychology, and Engineering. Conversely, scores are notably lower in Computer Science, History, and Physics, reflecting the inherent difficulty of handling highly technical content, specific historical contexts, and complex scientific reasoning in these fields. Overall, these results demonstrate that our evaluation framework effectively captures fine-grained and domain-specific performance, enabling a granular diagnostic analysis that goes beyond simple aggregate scoring to identify the varying strengths and weaknesses of agents across diverse academic and technical fields.
@@ 6.3 Correlation with Human Evaluation
To validate the proposed evaluation method and assess the contribution of each component, we compared LLM-based evaluations against human expert judgments. For each of the 45 reports, we collected two independent ratings from domain experts whose expertise aligned with the report’s topic (90 ratings total). We computed Pearson’s r, Spearman’s ρ, and LLM–human pairwise agreement for each of the five LLM-based evaluator models, and report the average of each metric across models. Additional details on the experimental setup and human evaluation protocol are provided in Appendix H.
@@ 6.4 Inter-evaluator Reliability
To verify that the proposed evaluation method yields consistent results across different LLM evaluators, we measure inter-evaluator reliability (Artstein, 2017; Lee et al., 2025). Specifically, we compute Krippendorff’s $\alpha$ and the intraclass correlation coefficient (ICC) using the scores from the five LLM judges described in Section 6.3.As shown in Table 3, Vanilla achieves moderate inter-evaluator reliability, but reliability drops substantially under +Dimensions. +Granular Rubrics yields a slight recovery, though still below Vanilla. +Expert Guidance attains the highest reliability across metrics. This suggests that expert guidance clarifies what to look for under each criterion, enabling more consistent judgments across different LLM judges.
@@ 6.5 Information Verification Module Evaluation
To evaluate the proposed Information Verification module, we present results on claim extraction and claim verification. These results quantify how our batch extraction and grouped verification design achieves strong performance while improving efficiency. Detailed results are provided in Appendix E and F.
@@ 7 Conclusion
Impact Statement
This paper presents methods for improving the efficiency and reliability of machine learning model evaluation and training. The proposed approach may support more systematic model comparison and selection, which can influence downstream deployment decisions. As with any automated evaluation framework, inappropriate or uncritical use may lead to over-reliance on quantitative metrics without sufficient human judgment. We therefore emphasize that the proposed methods are intended to complement, not replace, human oversight in model development and evaluation.
@@ Appendix A Comparison with Deep Research Benchmarks
In this section, we compare existing deep research benchmarks with our methodology. Table 6 extends and modifies the comparison table proposed in ResearchRubrics (Sharma et al., 2025), and adds new evaluation axes to summarize how DEER differs from prior work. The ResearchRubrics table uses five axes (whether rubrics are human-written, whether experts curate tasks, whether tasks are open-ended, whether non-technical domains are included, and whether an LLM-as-judge is used). On top of this, we add two axes that are essential for deep research report evaluation: (1) Claim fact-check and (2) Interpretability.
A.1 Benchmark Landscape
Search and browsing agent benchmarks.
@@ A.1 Benchmark Landscape
The first block of the table consists of benchmarks for evaluating search and browsing agents. AcademicBrowse (Zhou et al., 2025), BrowseComp (Wei et al., 2025a), and Mind2Web2 (Gou et al., 2025) evaluate, respectively, the ability to browse an academic corpus or the open web to produce short answers to complex queries, the ability to perform agentic search across diverse websites, and the consistency between generated answers and cited sources. These benchmarks focus on “how well the system finds information,” i.e., search and browsing strategies, and thus are one step removed from the deep research report quality evaluation that we study.
Benchmarks for partial abilities or conceptualizations of deep research.
@@
The second block covers benchmarks that focus on specific component abilities or conceptualizations of deep research. ExpertLongBench (Ruan et al., 2025) evaluates the ability to generate long-form expert text without external search. ResearchBench (Liu et al., 2025) evaluates the ability to extract inspirations from papers and generate research hypotheses. ResearcherBench (Xu et al., 2025) evaluates long-form responses to frontier AI research questions using a dual framework: expert-rubric insight quality and citation-based factuality (faithfulness/groundedness). ReportBench (Li et al., 2025a) and DeepScholar-Bench (Patel et al., 2025) assess academic survey/related-work reports, focusing primarily on literature selection and citation-grounded verifiability of report content. LiveDRBench (Java et al., 2025) evaluates the recovery of correct claims in search tasks that require many information units and non-trivial reasoning. SPOT (Son et al., 2025) measures how well a system can detect critical errors in published papers. These benchmarks finely evaluate partial abilities such as expert writing, hypothesis generation, claim discovery, and error detection, but it is difficult to view them as evaluating the overall quality of reports produced by web- or literature-based deep research agents.
Benchmarks for deep research report quality.
The third block targets benchmarks that evaluate the quality of deep research reports themselves. DeepResearchGym (Coelho et al., 2025) provides an offline web-corpus sandbox with an LLM-as-judge protocol. DeepResearchBench (Du et al., 2025) evaluates multi-domain web-based deep research reports using LLM-generated evaluation criteria (RACE), citation-based fact-checking (FACT), and dimensions including coverage, depth, presentation, and citation accuracy.
Concurrent work.
Concurrently with DEER, several benchmarks have further advanced report-level evaluation. DeepResearchArena (Wan et al., 2025) derives tasks from seminar transcripts and evaluates evidence–keypoint alignment (KSR/KCR/KOR) with task-specific checklists (ACE). RigorousBench (Yao et al., 2025) uses expert-curated queries and two-level human rubrics (GRR/QSR), and adds a trustworthiness signal via matching citations to curated trustworthy-source links (TSL). LiveResearchBench (Wang et al., 2025) evaluates web-based deep research reports in a live, multi-domain setting using LLM-judged criteria, e.g., presentation & organization, coverage & comprehensiveness, and citation accuracy. ResearchRubrics (Sharma et al., 2025) provides 2,500+ expert-written rubric items spanning axes such as explicit requirements, synthesis, and reference use, with mandatory vs. optional criteria per task.
A.2 Key Differences from Prior Work
Alignment with expert standards.
A key concern in deep-research evaluation is whether the reported score truly reflects expert notions of report quality. In several benchmarks (Du et al., 2025; Wang et al., 2025; Wan et al., 2025), LLMs are used not only as judges but also to instantiate parts of the evaluation criteria (e.g., LLM-generated dimensions or task-specific checklists), which can leave ambiguity about how closely evaluation aligns with expert standards; moreover, even with well-defined criteria, LLM judges may apply them unreliably due to limited domain knowledge and weak evidence judgment. In contrast, DEER anchors evaluation in a shared rubric system grounded in established expert reporting norms and guidelines, and further provides task-specific expert guidance so that LLM-based scoring better aligns with expert judgment rather than ad hoc scoring heuristics.
Claim-level fact checking.
Some benchmarks (Du et al., 2025; Wang et al., 2025) perform citation-based verification only for cited claims, leaving uncited claims unchecked. Another approach (Wan et al., 2025) scores alignment to keypoints extracted from cited URLs, making verification conditional on what is cited and less suited to detecting missing evidence. In contrast, DEER extracts all claims from a report and, for each claim, (i) determines whether evidence is required, (ii) links not only explicitly cited sources but also recovers omitted citation links by tracing each claim back to earlier cited context in the report, and (iii) verifies whether the linked evidence supports the claim. For more detailed comparisons, see Table 7.
Systematic interpretability.
Beyond a single overall score, deep-research evaluation should support consistent diagnosis of failure modes across tasks. When benchmarks use task- or prompt-specific sub-criteria, fine-grained diagnostics are not standardized across tasks, so results tend to be interpretable mainly at coarse, high-level dimensions and are difficult to compare at a shared checklist level (Du et al., 2025; Wang et al., 2025; Wan et al., 2025; Sharma et al., 2025). DEER instead uses a hierarchical, shared rubric taxonomy grounded in established expert report-writing norms and guidelines, applying a fixed set of rubric sub-dimensions and items across tasks. As a result, DEER evaluates each report against a dense set of rubric items for more thorough, fine-grained assessment, and supports rubric-item-level diagnosis of system weaknesses.
@@ B.1 Topic Domain Analysis
To construct deep research tasks, we analyzed 5,842 in-house user queries collected from our deep research system to estimate real-world domain demand and to guide the benchmark’s target domain distribution. Figure 4 shows the resulting distribution. Based on this analysis, we finalized 11 topic domains: the 10 most frequent domains and an Others category that aggregates all remaining domains. We used only aggregated topic counts/statistics derived from in-house queries; no raw queries were released or included, and all queries complied with the organization’s privacy policy, including the removal of any personally identifiable information.
@@ B.2 HLE Subject Mapping
Accordingly, we used Humanity’s Last Exam (HLE) (Phan et al., 2025) as a source of expert-written, high-difficulty seed items that align with expert-level report generation. To preserve the 11-topic domain composition derived from actual deep research logs, we mapped each domain to one of HLE’s 13 subject domains and sampled HLE items from the corresponding subjects as domain-specific seeds. At this time, we filtered candidates via a preliminary performance evaluation, excluding items that were excessively difficult for LLM-based evaluation and retaining only those within an appropriate difficulty range. As a result, we selected a total of 50 seed items across HLE’s 13 subject domains. The correspondence between deep research topic domains and HLE subject domains is summarized in Table 8.
@@ B.3 Conversion from HLE QA to Deep Research Reports
Each task query was drafted by one domain expert and cross-reviewed by another expert from the same field. Cross-review was repeated in multiple rounds as needed to check whether the reformulated task (i) is not overly narrow, (ii) requires expert-level domain expertise, (iii) specifies a sufficiently concrete research scope and direction, and (iv) does not contain excessive hints that could steer the model toward the answer or conclusion; the task was revised accordingly. Experts were individuals with a master’s degree in the relevant field or equivalent domain expertise. An example research task is provided in Figure 5.
@@ B.4 Construction of Expert Evaluation Guidance
For all 50 tasks, Expert Evaluation Guidance was written in the same expert workflow simultaneously with query reformulation, so that each task’s requirements and evaluation criteria are mutually aligned. As in query reformulation, each Guidance was drafted by one domain expert and cross-reviewed by another expert from the same field. During cross-review, we checked and revised whether (i) the topic’s mandatory content elements were included at a sufficiently detailed level and written in an evaluable way, (ii) optional content or stylistic preferences were not included as mandatory evaluation elements, and (iii) requirements specified in the task prompt’s writing requirements (writing-direction instructions/additional requested constraints) were reflected without omission in the Guidance. An example Expert Evaluation Guidance is provided in Figure 6.
@@ C.2 Construction Procedure and Evidence Mapping
Completeness aligns with guideline-level completeness checks that evaluate whether all required components of a report are present according to prescribed inclusion mandates across authoritative standards (Bastian and Moher, 2021; Boutron et al., 2010; International Organization for Standardization, 2011; What Works Clearinghouse, 2022; Percie du Sert et al., 2020; Global Reporting Initiative, 2023; von Elm and others, 2007). Scope reflects boundary-setting requirements that assess whether a report explicitly specifies its operative limits—such as eligibility conditions, assumptions, and constraints—within the structural fields defined in major evaluative frameworks (Bastian and Moher, 2021; National Institute of Standards and Technology, 2023; American Historical Association, 2024; American Economic Association, 2024; Center for Open Science, 2024; NASA, 2016; U.S. Securities and Exchange Commission, 2024; Chan and others, 2013). Helpfulness corresponds to coherence and utility standards that evaluate whether the report’s core conclusions not only align logically with the evidence base but also deliver sufficient specificity and practical feasibility to comprehensively address the user’s inquiry (Guyatt et al., 2013; American Psychological Association, 2025; Chang et al., 2024; Institute of Education Sciences, 2022; OECD, 2020).
@@
Quantification reflects evaluative requirements that check whether quantitative statements in a report correspond to verifiable computations, prespecified analytic procedures, and reproducible numerical evidence as established in major methodological frameworks (Guyatt et al., 2013; Joint Committee for Guides in Metrology, 2008; Association for Computing Machinery, 2025; International Union of Pure and Applied Chemistry, 2007; International Union of Pure and Applied Physics, 2010; The Econometric Society, 2024; American Educational Research Association et al., 2014; Mohr and others, 2024; OECD, 2011; International Accounting Standards Board, 2024; International Organization for Standardization, 2018; Intergovernmental Panel on Climate Change, 2019). Reasoning aligns with reasoning-assessment criteria that evaluate whether inferential steps are explicitly grounded in evidence, free of unsupported leaps, and consistent with theoretical or mathematical rigor (European Mathematical Society, 2025; American Philosophical Association, 2024; American Educational Research Association, 2006; American Mathematical Society, 2022; U.S. Congress, 2024; Tong et al., 2007).
@@
Introduction corresponds to structural-orientation requirements that assess whether a report’s opening section presents its scope and analytic pathway in accordance with prescribed organizational fields in established reporting and specification frameworks (Bastian and Moher, 2021; Boutron et al., 2010; IEEE Computer Society, 2018; IFRS Foundation, 2021; Gagnier and others, 2013). Body aligns with structural-progression criteria that evaluate whether the main sections develop the promised analytical sequence without omission or drift relative to the ordered components defined in recognized guideline structures (Bastian and Moher, 2021; Boutron et al., 2010; IEEE Computer Society, 2018; Eaton and others, 2024; Gagnier and others, 2013). Conclusion reflects closure-consistency requirements that examine whether final statements integrate preceding evidence without introducing unsupported expansions, consistent with the conclusion-governance rules embedded in major evaluative standards (Bastian and Moher, 2021; Boutron et al., 2010; Gagnier and others, 2013). Section-Level corresponds to intra- and inter-section coherence checks that assess alignment, ordering, and non-duplication of content in accordance with structured specification and reporting templates in authoritative frameworks (Bastian and Moher, 2021; IEEE Computer Society, 2018; IFRS Foundation, 2021).
@@
Report Format corresponds to format-governance requirements that evaluate whether a report adheres to prescribed structural conventions for scientific communication as codified in established editorial and reporting standards (Bastian and Moher, 2021; Boutron et al., 2010; International Committee of Medical Journal Editors, 2025; Comrie, B. and Haspelmath, M. and Bickel, B., 2015; Society for Industrial and Applied Mathematics, 2024; International Phonetic Association, 1999; American Political Science Association, 2018; American Sociological Association, 2022). Writing Quality aligns with language-precision criteria that assess clarity, specificity, and terminological consistency according to recognized guidelines for accurate and unbiased scholarly expression (Boutron et al., 2010; International Committee of Medical Journal Editors, 2025; University of Chicago Press, 2024; Garner, 2019, 2013). Paragraph Quality reflects cohesion-assessment rules that examine whether individual paragraphs follow a coherent internal logic—anchoring topic statements, supporting evidence, and transitional structure—in line with authoritative reporting and specification frameworks (Bastian and Moher, 2021; IEEE Computer Society, 2018; Harvard Law Review Association, 2020; American Sociological Association, 2022). Readability corresponds to accessibility-oriented requirements that evaluate whether the narrative facilitates comprehension through appropriate signaling, explanatory devices, and presentation practices as outlined in major editorial and reporting guidelines (Bastian and Moher, 2021; International Committee of Medical Journal Editors, 2025; Garner, 2013; Global Reporting Initiative, 2023).